OpenStudy (anonymous):

I don't understand the point of cofunction identities, like I know how to solve problems with it but I don't know how to make sense of it. Can someone please explain it?

OpenStudy (anonymous):

\[f(\frac{\pi}{2}-\theta)=cof(\theta)\]

OpenStudy (anonymous):

for example \[\sin(\frac{\pi}{2}-\theta)=\cos(\theta)\] and so on